Paul has $70,000 to invest. His intent is to earn 15% interest on his investment. He can invest part of his money at 8% interest and part at 18% interest. How much does Paul need to invest in each option to make a total 15% return on his $70?
Found 2 solutions by solver91311, greenestamps:
Answer by solver91311(24713)





Solve for and then calculate


John

My calculator said it, I believe it, that settles it


Answer by greenestamps(13198)



The setup for solving the problem shown by the other tutor is a good example of how to solve a problem using formal algebra. You should understand how to use that method.

If a formal algebraic method is not required, here is a solution method that is almost trivial.

The target return rate, 15%, is 7/10 of the way from 8% to 18%. (8 to 15 is a difference of 7; 8 to 18 is a difference of 10.)

Therefore, 7/10 of the total needs to be invested at the higher rate.

ANSWER: 7/10 of the $70,000, or 49,000, at 18%; the other $21,000 at 8%.

CHECK:
0.18(49000)+0.08(21000) = 8820+1680 = 10500
0.15(70000) = 10500